### 18 Dec 2023 v2

# Learning from History: Task-agnostic Model Contrastive Learning for Image Restoration

Abstract
Contrastive learning has emerged as a prevailing paradigm for high-level vision tasks, which, by introducing properly negative samples, has also been exploited for low-level vision tasks to achieve a compact optimization space to account for their ill-posed nature. However, existing methods rely on manually predefined and task-oriented negatives, which often exhibit pronounced task-specific biases. To address this challenge, our paper introduces an innovative method termed ’learning from history’, which dynamically generates negative samples from the target model itself. Our approach, named Model Contrastive paradigm for Image Restoration (MCIR), rejuvenates latency models as negative models, making it compatible with diverse image restoration tasks. We propose the Self-Prior guided Negative loss (SPN) to enable it. This approach significantly enhances existing models when retrained with the proposed model contrastive paradigm. The results show significant improvements in image restoration across various tasks and architectures. For example, models retrained with SPN outperform the original FFANet and DehazeFormer by 3.41 dB and 0.57 dB on the RESIDE indoor dataset for image dehazing. Similarly, they achieve notable improvements of 0.47 dB on SPA-Data over IDT for image deraining and 0.12 dB on Manga109 for a 4$\times$ scale super-resolution over lightweight SwinIR, respectively. Code and retrained models are available at https://github.com/Aitical/MCIR.
Introduction
Image restoration, aiming at recovering a high-quality image from the degraded one, is a fundamental problem in the fields of image processing and computer vision (Yang et al. 2021; Wang, Chen, and Hoi 2021a; Jiang et al. 2023). Deep learning approaches have brought about considerable advancements in image restoration while there are still challenges due to its ill-posed nature (Wang, Chen, and Hoi 2021b; Lu et al. 2023). The success of self-supervised learning paradigm for high-level tasks, especially those contrastive learning methods, has drawn great attention (Gui et al. 2023a). This inspires many researchers to made strides in improving the end-to-end learning paradigm for image restoration tasks, incorporating the concept of sample contrastive learning and bridging the gap between high-level and low-level tasks (Wu et al. 2021; Chen et al. 2022b; Ye et al. 2022; Wu, Jiang, and Liu 2023; Zheng et al. 2023). Image restoration tasks usually contain high-quality ground truth as learning target (positive sample), and more attention is paid on mining appropriate negative examples. For instance, Wu et al. (Wu et al. 2021) directly utilized the low-quality input as the negative sample and introduced the contrastive paradigm for the image dehazing task. Wu et al. proposed a hard negative construction for image super-resolution tasks. Compared to the super-resolved results, hard negatives with similar image quality to the anchor sample can push it ahead effectively (Wu, Jiang, and Liu 2023). Most recently, Zheng et al. proposed a progressively improved negative lower bound for image dehazing, employing multiple pre-trained models to provide consensual negatives (Zheng et al. 2023).
Figure 1: Illustration of the proposed model contrastive paradigm. We provide a common optimization space for it. To the target model $f_{\theta}$, the proposed model contrastive paradigm exploits negative samples from the latency model $f_{\theta^{{}^{\prime}}}$ smoothly. Compared to task-oriented negatives in previous work, our model contrastive paradigm is task-agonist and general to various image restoration tasks. This provides a compact optimization space adaptively (pushing target model $f_{\theta}$ closer to assumed optimal $f_{\theta^{*}}$).
While existing contrastive learning methods for low-level tasks, leveraging potent approaches like hard negative mining (Wu, Jiang, and Liu 2023) and curriculum learning strategies (Zheng et al. 2023), have shown impressive performance, certain intrinsic limitations persist. Chiefly, there exists an over-reliance on task-oriented prior, leading to limited generalization capability across multiple image restoration tasks. Many existing methods exhibit a pronounced task-oriented bias, where negative sample generation is often influenced by prior knowledge and empirical evaluation centered around a singular target task (Wu, Jiang, and Liu 2023; Zheng et al. 2023). Such approaches, while effective in isolation, often impede to a diverse set of image restoration tasks and model architectures. In light of these challenges, it compels us to ask: Is there a task-agnostic and general method for negative sampling that could potentially enhance the performance of a diverse range of image restoration tasks?
Taking this into consideration, we turn our focus to the target model itself, rather than sample selection. In recent literature, much of the spotlight has been directed towards negative sample collection, often overlooking a latent gem—the latency model—during the learning process. Specifically, the latency model, when operating within a close optimization step, shares strikingly similar parameters with the current model. This intrinsic similarity paves the way for the construction of proper negatives pertinent to the current anchor sample. To illustrate this concept, we introduce a toy example in Fig. 1. It becomes evident that, throughout the learning journey, the output of the latency model exhibits a sub-optimal but congruent distribution relative to the current anchor sample. This alignment offers the potential to derive ’hard’ negatives that are well-suited to the task at hand. Furthermore, as the entire learning process is incrementally refined, the negatives in our model contrastive paradigm adopt a curriculum way naturally. Motivated by these insights, we put forth an innovative model contrastive paradigm for image restoration tasks. In contrast to the previous approaches with complex negative mining strategies, the core of our proposed model contrastive paradigm lies in the self-promoted negatives between latency and current models. Notably, it is task-agonist and general to diverse image restoration tasks.
In detail, we proposed a model contrastive paradigm for various image restoration tasks (MCIR) with a Self-Prior guided Negative loss (SPN). Importantly, the proposed MCIR is compatible with existing approaches across different architectures and tasks. We retrain existing models by the proposed model contrastive and some results are presented in Fig. 2. One can find that retrained models achieve promising improvement across multiple tasks and architectures. More specifically, based on the lightweight EDSR (Lim et al. 2017) and SwiniR (Liang et al. 2021), our retrained models display superior performance on Manga109 test dataset for ×$\times 4$ scale image super resolution, boasting improvements of 0.16 dB and 0.12 dB in term of PSNR. Retrained IDT (Xiao et al. 2022) achieves 0.38 dB and 0.7 dB gains on Rain200L and SPA datasets for image deraining. Moreover, for image dehazing, we gain a notable improvement of 3.41 dB and 0.57 dB compared to the original FFANet (Qin et al. 2020a) and DehazeFormer (Song et al. 2023), respectively.
The main contributions of this work are:
  • The paper proposes a novel approach MCIR through a task-agnostic model contrastive paradigm, which provides the adaptive generation of negative samples directly from the target model itself. Unlike conventional methods that manually apply negative samples to specific target task, the proposed model contrastive paradigm exhibits versatility across multiple tasks and models.
  • The paper introduces the Self-Prior guided Negative loss (SPN) for image restoration, seamlessly compatible with existing methods. SPN provides a simple to enhance existing image restoration models by integrating self-supervision principles within our model contrastive paradigm.
  • The paper demonstrates the effectiveness of the proposed approach by retraining existing models with the MCIR, which significantly improves image restoration across various tasks and architectures.
Figure 2: Comparisons between models retrained by our proposed model contrastive paradigm and the originals. Retrained models can achieve remarkable improvements on various image restoration tasks. Details are provided in Section Experiment
Related Work
Image Restoration
Image restoration, aiming at recovering a high-quality image from the degraded one, is a fundamental problem in the fields of image processing and computer vision. It is indispensable in several domains, including photography, surveillance, medical imaging, and remote sensing (Su, Xu, and Yin 2022). In recent years, approaches based on deep learning have emerged as potent tools, providing significant advances for image restoration (Wang, Chen, and Hoi 2021a; Jiang et al. 2023; Wang et al. 2021a; Gui et al. 2023b). Among these, Convolutional Neural Networks (CNN) have been utilized, resulting in a plethora of methods designed to address various image restoration tasks (Dong et al. 2016; Zhang et al. 2017; Zhang, Zuo, and Zhang 2018; Yang et al. 2017a). Furthermore, remarkable strides have been made by the development of a spectrum of deep models, including proposals for residual or dense architectures (Lim et al. 2017), the integration of attention mechanisms (Zhang et al. 2018; Liu et al. 2019), and the implementation of UNet-based architectures (Ronneberger, Fischer, and Brox 2015; Jiang et al. 2020). In addition, some work explored the attention mechanism to boost the performance for image super-resolution (Zhang et al. 2018; Liu et al. 2018; Mei, Fan, and Zhou 2021), image deraining (Li et al. 2018a; Wang et al. 2019a), image deblurring (Suin, Purohit, and Rajagopalan 2020), and so on. On the frontier of recent development, Transformer architectures achieves significant breakthroughs in computer vision (Liu et al. 2021; Dosovitskiy et al. 2021). This trend has spurred the development of Transformer-based interventions for image restoration tasks such as image super-resolution (Liang et al. 2021), image dehazing (Song et al. 2023; Zamir et al. 2022), image deraining (Xiao et al. 2022), image deblurring (Wang et al. 2022), and low-light image enhancement (Wang et al. 2023). We are witnessing a surge in the performance of increasingly complex architectures and there is a growing demand for applications in real-world systems. Many researcher have turned attention to lightweight methods (Jiang et al. 2021; Lamba and Mitra 2021; Hui et al. 2019; Wu et al. 2023b; Lu et al. 2022; Wu et al. 2023a).
Contrastive Learning
Self-supervised learning has achieved great success in high-level vision tasks, especially those contrastive learning methods (Gui et al. 2023a). For low-level image restoration tasks, many researchers attempted to bridge the gap between low-level and high-level tasks and exploit more self-supervised information to promote image restoration tasks (Wang et al. 2021b; Wu et al. 2021; Chen et al. 2022b; Ye et al. 2022; Zheng et al. 2023; Wu, Jiang, and Liu 2023). In contrast to unsupervised learning for high-level vision tasks, low-level tasks usually have high-quality ground-truth images. Discovering informative negative samples is crucial for low-level tasks. Wu et al. took the low-quality input as the negative sample and proposed a perceptual-based contrastive loss to achieve a more compact convergence space (Wu et al. 2021). Wu et al. proposed a practical contrastive learning framework for super-resolution single image and proposed the hard negative construction (Wu, Jiang, and Liu 2023). By interpolating the negative information, they further improve the performance for target image restoration task. However, adaptation and generalization of these methods to various image restoration models and tasks are limited.
Compared to existing methods (Wu et al. 2021; Wu, Jiang, and Liu 2023; Zheng et al. 2023), our work presents a task-agnostic model contrastive paradigm for image restoration, utilizing a self-prior guided negative loss. This approach simplifies negative sample generation and offers a versatile framework for various image restoration tasks, setting it apart from existing methods.
Method
Figure 3: Comparisons of different learning paradigms for image restoration tasks. (a) Depicts the vanilla end-to-end learning paradigm, where target image restoration models are trained with supervised reconstruction loss; (b) Illustrates the sample contrastive paradigm adopted in existing methods, leveraging additional information from manually defined negative samples; (c) Showcases our model contrastive paradigm, which generates negative samples from the latency target model, providing a task-agnostic and adaptive approach to various image restoration tasks.
Overall Framework
Deep learning-based image restoration methodologies have recently made considerable breakthroughs across multiple restoration tasks, including image super-resolution, image dehazing, image deblurring, image deraining, and so on. In this paper, we focus our attention on the vanilla end-to-end learning process in deep learning methods. To facilitate clarity, we provide a succinct illustration of the end-to-end image restoration framework, visualized in Fig. 3 (a). Given a low-quality input image denoted as $I^{LQ}$, the target model $f_{\theta}$, where $\theta$ represents model parameters, processes low-quality input to produce a reconstructed image, $I^{Rec}$. Concurrently, its high-quality counterpart is denoted by $I^{HQ}$ and the optimizing of the target model is steered by the minimization of the reconstruction loss $\mathcal{L}_{rec}$. Delving further, Fig. 3 (b) showcases the contrastive learning strategies tailored for image restoration tasks. Typically, researchers harness the power of negative samples, exploiting informative task-prior specific negatives and extend the vanilla end-to-end learning paradigm by the negative constraint $\mathcal{L}_{neg}$ (Wu et al. 2021; Zheng et al. 2023; Wu, Jiang, and Liu 2023).
In this paper, we propose an innovative model contrastive paradigm for image restoration tasks, shown in Fig. 3 (c). Instead of utilizing task-specific and predefined negative samples, we craft our negatives directly from the target model itself. At this point, we denote the target model $f_{\theta}$, indicative of the optimized one at the current training iteration $t$. Subsequently, we employ a distinct step, $s$, to establish a latency model $f_{\theta^{{}^{\prime}}}$, also referred to as the negative model, to generate adaptive negative samples for the training of the target model. The overall pipeline is straightforward, and the detailed implementation of negative models and associated loss functions will be discussed in the following sections.
Learning from History
We propose a practical implementation for leveraging latency models to generate informative negative samples during training, as introduced earlier. This method circumvents the challenges of significant gaps between model checkpoints and the impracticality of frequent processing, especially with larger models. In detail, we introduce the efficient strategy of exponential moving averages (EMA) to achieve a smooth negative model. The update equation for the negative model is:
$\begin{split}&\theta^{{}^{\prime}}=w\theta^{{}^{\prime}}+(1-w)\theta,\\ &\mathrm{s.t.}~{}t\%s=0.\end{split}$ (1)
In this schema, $w$ denotes the update weight, $s$ is the update step, and $t$ captures the current iteration. To ensure the preservation of latency parameters, we adopt a long step $s$ and selectively update the negative model $f_{\theta}$ at intervals of every $s$ iterations.
Self-Prior Guided Negative Loss
At the heart of our model contrastive paradigm, a loss function mediates between the target reconstruction $I^{Rec}$ and its negative counterpart, $I^{Neg}$. Herein, we take the pre-trained VGG (Simonyan and Zisserman 2015) network as the embedding network to map the samples into a latent feature space, where $f^{Rec}=\text{VGG}(I^{Rec})$ and $f^{Neg}=\text{VGG}(I^{Neg})$. Then the proposed negative loss $\mathcal{L}_{neg}$ is formulated as
$\mathcal{L}_{neg}=\|f^{Rec}-f^{Neg}\|_{1}.$ (2)
Furthermore, our model-based contrastive paradigm can incorporate multiple negatives by adding more negative models. For more robustness, we take multiple distinct steps to obtain several latency models. The combined negative loss, accounting for multiple negatives, is represented as
$\mathcal{L}^{N}_{neg}=\frac{1}{N}\sum_{i=1}^{N}{\|f^{Rec}-f^{Neg}_{i}\|}_{1},$ (3)
where $N$ denotes the total number of negative models and $f_{i}^{Neg}$ corresponds to the latent feature of the $i$th negative sample.
Compared to existing methods, our negative regularization is a self-prior guided loss function, where the negative samples stem from the target model itself. More important, it is general and transferable to existing image restoration models while retaining the original learning strategy. Typically, image restoration tasks utilize a reconstruction loss, $\mathcal{L}{rec}$, relying on metrics such as Mean Absolute Error (MAE) and Mean Squared Error (MSE). To validate the prowess of our model-based contrastive paradigm by $\mathcal{L}_{neg}$, we have retrained numerous exiting methods with it, testing across different image restoration tasks and architectures. The precise formulation of the reconstruction loss $\mathcal{L}_{rec}$ is contingent on the method under retraining.
Generally, the total loss function within our model-based contrastive paradigm is defined as:
$\mathcal{L}=\mathcal{L}_{rec}-\lambda\mathcal{L}^{N}_{neg},$ (4)
where $\mathcal{L}_{rec}$ represents the corresponding reconstruction loss adopted in existing method, and $\lambda$ is the balancing coefficient. This simple formulation allows us to incorporate the proposed SPN with existing image restoration methods, enhancing their flexibility and adaptability to various tasks.
Method Architecture Scale
Avg.
PSNR/SSIM
Set14
PSNR/SSIM
B100
PSNR/SSIM
Urban100
PSNR/SSIM
Manga109
PSNR/SSIM
EDSR-light CNN ×$\times 2$ 32.06/0.9303 33.57/0.9175 32.16/0.8994 31.98/0.9272 30.54/0.9769
+SPN (Ours) 32.19/0.9313 33.67/0.9182 32.21/0.9001 32.23/0.9297 30.64/0.9772
EDSR-light ×$\times 4$ 28.14/0.8021 28.58/0.7813 27.57/0.7357 26.04/0.7849 30.35/0.9067
+SPN (Ours) 28.21/0.8040 28.63/0.7829 27.59/0.7369 26.12/0.7878 30.51/0.9085
SwinIR-light Transformer ×$\times 4$ 28.46/0.8099 28.77/0.7858 27.69/0.7406 26.47/0.7980 30.92/0.9151
+SPN (Ours) 28.55/0.8114 28.85/0.7874 27.72/0.7414 26.57/0.8010 31.04/0.9158
SwinIR ×$\times 4$ 28.88/0.8190 28.94/0.7914 27.83/0.7459 27.07/0.8164 31.67/0.9226
+SPN (Ours) 28.93/0.8198 29.01/0.7923 27.85/0.7465 27.14/0.8176 31.75/0.9229
Table 1: Comparison results of image super-resolution. We take CNN-based EDSR (Lim et al. 2017) and Transformer-based SwinIR (Liang et al. 2021) as our baselines. Results of retrained models by the proposed model contrastive paradigm are in bold. Avg. presents the mean value of four test datasets.
Figure 4: Visual comparison for image super-resolution tasks. Displayed are results from both the original SwinIR models and those retrained by our model contrastive paradigm. The enhancements brought about by our approach are clearly evident.
Remark
In this study, we propose the model contrastive paradigm for image restoration tasks. This approach significantly simplifies the construction of negatives while simultaneously providing adaptive and effective ones. Contrary to existing approaches (Wu et al. 2021; Wu, Jiang, and Liu 2023; Zheng et al. 2023), the proposed model contrastive paradigm does not rely on the restrictions of task-oriented priors, making it versatile and universal to various image restoration tasks. In essence, our contribution lies in providing a task-agonist and general approach to the construction of negative samples, expanding the scope of contrastive learning in the field of image restoration.
Experiment
Experimental Settings
Image Super-Resolution
We retrain the CNN-based EDSR (Lim et al. 2017) and Transformer-based SwinIR (Liang et al. 2021). We take 800 images from DIV2K (Agustsson and Timofte 2017) for training and test them on five benchmark datasets.
Image Dehazing
We employ the CNN-based FFANet (Qin et al. 2020a) and Transformer-based DehazeFormer (Song et al. 2023) as our baselines and retrain them for image dehazing tasks. Following (Song et al. 2023), we utilize the indoor training dataset (ITS) and RESIDE-6K to separately train the models and evaluate on the synthetic objective testing set (SOTS).
Image Deblurring
For image deblurring, we retrain and evaluate NAFNet (Chen et al. 2022a) on GoPro dataset.
Image Deraining
For image deraining, we conduct experiments on several publicly available benchmarks, namely Rain200L/H (Yang et al. 2017b), DID-Data (Li et al. 2018b), DDN-Data (Fu et al. 2017), and SPA-Data (Wang et al. 2019b). We adopt the Transformer-based IDT model (Xiao et al. 2022) as our baseline and subsequently retrain it using the proposed model contrastive paradigm.
Given the diverse training settings across various image restoration tasks, we ensure a fair comparison by strictly adhering to the original training configurations of each retrained model, as specified in their literature.
Comparison Results
Method
Avg.
PSNR/SSIM
Rain100L
PSNR/SSIM
Rain100H
PSNR/SSIM
DID
PSNR/SSIM
DDN
PSNR/SSIM
SPA
PSNR/SSIM
(CVPR21) MPRNet 36.17/0.9543 39.47/0.9825 30.67/0.9110 33.99/0.9590 33.10/0.9347 43.64/0.9844
(AAAI’21) DualGCN 36.69/0.9604 40.73/0.9886 31.15/0.9125 34.37/0.9620 33.01/0.9489 44.18/0.9902
(ICCV’21) SPDNet 36.54/0.9594 40.50/0.9875 31.28/0.9207 34.57/0.9560 33.15/0.9457 43.20/0.9871
(CVPR’22) Uformer-S 36.95/0.9505 40.20/0.9860 30.80/0.9105 34.46/0.9333 33.14/0.9312 46.13/0.9913
(CVPR’22) Restormer 37.49/0.9530 40.58/0.9872 31.39/0.9164 35.20/0.9363 34.04/0.9340 46.25/0.9911
(CVPR’23) DRSformer 38.33/0.9676 41.23/0.9894 32.17/0.9326 35.35/0.9646 34.35/0.9588 48.54/0.9924
(TPAMI’23) IDT 37.77/0.9593 40.74/0.9884 32.10/0.9343 34.85/0.9401 33.80/0.9407 47.34/0.9929
(Ours) IDT+SPN 38.03/0.9610 41.12/0.9893 32.17/0.9352 34.94/0.9424 33.90/0.9442 48.04/0.9938
Table 2: Comparison results of image deraining. We take IDT (Xiao et al. 2022) as the benchmark and retrain it with the proposed model contrastive paradigm. We evaluate the performance on several image deraining datasets, and our results are in bold (Average performance of the five datasets is calculated in the Avg. column).
Figure 5: Visual comparisons between IDT and our retrained one. A divergence map delineates the differences between the IDT output and ours, highlighting the improvement, particularly in degraded regions.
Image Super-Resolution
We first investigate the image super-resolution task, utilizing both CNN-based EDSR and Transformer-based SwinIR as baselines to evaluate our proposed self-pprior guided negative regularization. Results presented in Tab. 1 across various architectures and model capacities. Our proposed model contrastive paradigm can achieve considerable improvements when retraining lightweight models such as EDSR and SwinIR, achieving average gains of 0.07 dB and 0.09 dB on scale $\times$4, respectively. Moreover, the retrained EDSR model demonstrats a noteworthy average improvement of 1.3 dB for the scale $\times$2 task. Notably, there are substantial improvements by our retrained on the Urban100 and Manga109 datasets. This implies that our model contrastive paradigm facilitates a more accurate and compact convergence in the restoration process. Moreover, comparisons with current state-of-the-art methods are presented in Fig. 2 (a). Remarkably, the retrained EDSR and SwinIR can eclipse the performance of several contemporary methods. These underscore the promise of our model contrastive paradigm.
In addition to the lightweight models, we also retrain the large SwinIR model. Our proposed approach exhibits its general applicability to larger models, providing an average improvement of 0.05 dB. It is worth noting that the improvements seen in larger models were relatively smaller when compared to the lightweight models. This observation is consistent with the understanding that lightweight models, owing to their smaller capacity, may struggle to find an optimal solution. Therefore, our proposed model contrastive paradigm provides significant assistance in these cases. Conversely, larger models have the capacity to fit or even overfit the dataset, hence the model contrastive paradigm improves them within a lesser margin. Visual results are showcased in Fig. 4. It is also evident that our model contrastive paradigm can improve existing methods, particularly in bringing more precise textures.
Image Deraining
We take the state-of-the-art IDT (Xiao et al. 2022) as the baseline and retrain it by the proposed model contrastive paradigm on five datasets ,and Tab. 2 provides a comprehensive comparison. The retrained IDT, achieves an average PSNR improvement of 0.26 dB across all test datasets. For instance, it showcase a gain of 0.7 dB in comparison to its original on SPA dataset. Furthermore, visual results are illustrated in Figure 5, accompanied by a divergence map. It intuitively highlights the enhancements in the degraded regions. These insights underscore the capability of our model contrastive paradigm to refine the target degradation.
Methods ITS RESIDE-6K
SOTS-indoor SOTS-mix
PSNR SSIM PSNR SSIM
(CVPR’21) AECR-Net 37.17 0.990 - -
(ICLR’23)  SFNet 41.24 0.996 - -
[…]
(AAAI’20) FFANet 36.39 0.989 29.96 0.973
(Ours) FFANet+SPN 39.80 0.995 30.65 0.976
(TIP’23) DehazeFormer-T 35.15 0.989 30.36 0.973
(Ours) DehazeFormer-T+SPN 35.51 0.990 30.44 0.974
(TIP’23) DehazeFormer-S 36.82 0.992 30.62 0.976
(Ours) DehazeFormer-S+SPN 37.24 0.993 30.77 0.978
(TIP’23) DehazeFormer-B 37.84 0.994 31.45 0.980
(Ours) DehazeFormer-B+SPN 38.41 0.994 31.57 0.981
Table 3: Comparison results of image dehazing. Reults of our retrained models are in bold.
Figure 6: Visual results of FFANet and our retrained one for image dehazing.
Method MIMO-UNet HINet MAXIM Restormer UFormer NAFNet NAFNet+SPN (Ours)
PSNR 32.68 32.71 32.86 32.92 32.97 32.87 32.93 (+0.06)
SSIM 0.959 0.959 0.961 0.961 0.967 0.9606 0.9619 (+0.0013)
Table 4: Comparison results of image deblurring. We take NAFNet (Chen et al. 2022a) as the benchmark and retrain it with the proposed model contrastive paradigm on GoPro dataset.
Image Dehazing
Following the recent work (Song et al. 2023), we utilize the SOTS-indoor and SOTS-mix datasets for testing. In detail, we retrain FFANet (Qin et al. 2020b) and DehazeFormer (Song et al. 2023), with the outcomes detailed in Fig. 3. From the table, a remarkable enhancement is achieved over FFANet, exhibiting gains of 3.41 dB and 0.69 dB on the indoor and mixed test datasets, respectively. Concurrently, the Transformer-based model IDT also record advancements, evident across various model scales. Specifically, the retrained DehazeFormer-B achieves the most pronounced improvement of 0.57 dB on the indoor test dataset. Intriguingly, the retrained FFANet even outperformes the DehazeFormer. It is reasonable that FFANet is a CNN-based model, which has a large model capacity with heavy parameter counts, and our model contrastive paradigm appears to steer FFANet towards more optimized results. Some visual results are presented in Fig. 6. One can find that retrained FFANet is clearer with less artifacts.
Image Debluring
Results of the retrained NAFNet (Chen et al. 2022a) for image deblurring are in Tab. 4. One can find that the retrained NAFNet showcases notable enhancements compared to the original and outperforms the Transformer-based Restormer (Zamir et al. 2022).
Methods Task & Dataset PSNR SSIM
FFANet (Qin et al. 2020b) 36.39 0.9886
+CR (Wu et al. 2021) Image Dehazing 36.74 0.9906
+CCR (Zheng et al. 2023) (SOTS-indoor ) 39.24 0.9937
+SPN (Ours) 39.80 0.9947
EDSR (Lim et al. 2017) SISR 26.04 0.7849
+PCL (Wu, Jiang, and Liu 2023) (Urban100) 26.07 0.7863
+SPN (Ours) 26.12 0.7878
Table 5: A comparative analysis of the existing contrastive paradigms versus our proposed model contrastive approach. Existing methods are task-oriented and proposed for image dehazing and image super-resolution separately. Our model contrastive paradigm is task-agonist and general to improve diverse image restoration tasks.
Comparison to Existing Sample Contrastive Paradigm
In contrast to existing methods that employ task-oriented negatives, our model contrastive paradigm provides straightforward enhancement to image restoration tasks. We furnish a comprehensive comparison between exiting contrastive approaches and our innovative model contrastive paradigm. For image dehazing task, we examine previous methods based on FFANet, including contrastive regularization (CR) (Wu et al. 2021), curricular contrastive regularization (CCR) (Zheng et al. 2023), and our self-prior guided negative loss (SPN). The comparative results can be found in Tab. 5. Instead of merely employing low-quality input images as negative samples, CCR achieves significant advancements by utilizing consensus negatives from pre-trained models and the curriculum learning approach. Intriguingly, our SPN outperforms other methods, achieving the best improvement without any additional prior knowledge. In addition, for image super-resolution (Wu, Jiang, and Liu 2023), our model contrastive paradigm again registers the highest gains, as presented in Tab. 5. This consistent performance underscores the fact that our proposed model contrastive paradigm is task-agnostic and highly effective for diverse image restoration tasks.
Negative Model - Random Pre-trained Default
Avg. PSNR 28.46 28.44 28.51 28.55
Table 6: Results of retrained lightweight SwinIR with different negative models.
Ablation Studies
Impact of Negative Model
To provide a deeper understanding for the role of our negative model, our ablation study involves the exploration of different configurations of the negative model, including an randomly initialized and fixed SwinIR model, a pre-trained and fixed SwinIR model, and our default one. The results of these different configurations, which shed light on the efficacy of our proposed approach, are presented in Tab. 6. ’-’ presents the baseline model without contrastive negative model. Firstly, negative model with random and fixed parameters can not afford effective negatives and impedes the target SR model. When we set the negative model as the pre-trained and fixed one, it enhances the target model but achieves worse performance than ours. We think the adaptive and curricular negatives generated by model contrastive paradigm is essential to provide a smooth negative regularization, which is consistent with some existing methods.
Step $s$ 100 500 1000 2000 All
FFANet 38.90 38.54 38.27 37.55 39.80
Table 7: Ablation studies of the negative step $s$. We retrain FFANet with different negative step $s$ on indoor dataset.
Impact of Negative Step
We conducted an ablation study, where FFANet was retrained with various negative models, each differentiated by its updating iteration, $s$. The results are tabulated in Tab. 7. It becomes evident that the choice of step $s$ distinctly influences both the negative model and performance. A smaller step provides more challenging negatives compared to the larger step, leading to superior performance. In our experiments, we employ multiple negative models with the four steps together, achieving the best results. The incorporation of multiple negative models ensures a consistent supply of robust negatives.
Impact of Balancing Coefficient
We study the influence of the coefficient $\lambda$ in Eq. (1), and results are presented in Tab. 8, where 0 value means the original $L_{1}$ loss without our negative regularization. When $\lambda$ set as 1e-2 or 1e-3, there is collapse during training, especially in the early stage. This is reasonable that large negative loss can influence the model updating towards to the optimal direction. When $\lambda$ is 1e-4, our model contrastive paradigm achieves the best performance. Considering that the proposed model contrastive paradigm is task-agonist and adaptive to various image restoration tasks, in this paper, we take $\lambda=1e-4$ as the default value across different tasks.
$\lambda$ Value 0 1e-2 1e-3 5e-4 1e-4 1e-5
EDSR-light 28.58 - - 28.56 28.63 28.60
Table 8: Ablation studies on coefficient in total loss function.
$w$ in EMA 0 0.01 0.1 0.5 0.9 0.999
EDSR-light - - 28.63 28.60 26.60 28.61
Table 9: Ablation studies on the updating weight in EMA.
Impact of Updating Weight in EMA
In reference to Eq. (1), we examine the influence of varying updating weights. The outcomes are detailed in Tab. 9. In general, a smaller value of $w$ retains less latency information, bringing it in closer alignment with the target model. This results in the generation of the most challenging negatives. The symbol ’-’ indicates model collapse during training. As $w$ increases, it becomes evident that the optimal performance is attained at $w=0.1$. Based on our empirical observations, we consistently utilize $w=0.1$ as the default setting in all subsequent experiments.
Discussion and Limitation
In this paper, we proposed a task-agnostic model contrastive paradigm for low-level tasks, improving existing models. While our experiments cover a range of tasks, areas like JPEG artifact removal, image denoising, and certain real-world scenarios remain unexplored. Our ablation studies focused on image super-resolution, as shown in Tables 6 and 7, but we did not extensively evaluate our hyper-parameters across all tasks. These unexamined aspects offer valuable opportunities for future research.
Conclusion
In this study, we propose innovative model contrastive paradigm for various low-level tasks. Compared to the task-oriented negatives in existing methods, the proposed model contrastive paradigm, constructing negatives from the target model itself, a is task-agonist, general to various image restoration tasks by a Self-Prior guided Negative loss (SPN). Our proposed SPN is straightforward to implement and does not rely on any task or data prior. We have retrained several image restoration models, and they achieve significant improvements across various tasks and architectures. In the future, we believe it would be meaningful to evaluate our proposed paradigm in more dense prediction tasks, potentially offering fresh insights and advances for the community.
Acknowledgments The research was supported by the National Natural Science Foundation of China (U23B2009, 92270116).

### 19 Jan 2024 v4

@@
 The results show significant improvements in image restoration across various tasks and architectures.
-For example, models retrained with SPN outperform the original FFANet and DehazeFormer by 3.41 dB and 0.57 dB on the RESIDE indoor dataset for image dehazing.
+For example, models retrained with SPN outperform the original FFANet and DehazeFormer by 3.41 and 0.57 dB on the RESIDE indoor dataset for image dehazing.
 Similarly, they achieve notable improvements of 0.47 dB on SPA-Data over IDT for image deraining and 0.12 dB on Manga109 for a 4$\times$ scale super-resolution over lightweight SwinIR, respectively.
@@
 2023).
-Deep learning approaches have brought about considerable advancements in image restoration while there are still challenges due to its ill-posed nature (Wang, Chen, and Hoi 2021b; Lu et al.
+Deep learning approaches have made considerable advancements in image restoration, while there are still challenges due to its ill-posed nature (Wang, Chen, and Hoi 2021b; Lu et al.
 2023).
-The success of self-supervised learning paradigm for high-level tasks, especially those contrastive learning methods, has drawn great attention (Gui et al.
+The success of the self-supervised learning paradigm for high-level tasks, especially those using contrastive learning methods, has drawn great attention (Gui et al.
 2023a).
-This inspires many researchers to made strides in improving the end-to-end learning paradigm for image restoration tasks, incorporating the concept of sample contrastive learning and bridging the gap between high-level and low-level tasks (Wu et al.
+This inspires many researchers to make strides in improving the end-to-end learning paradigm for image restoration tasks, incorporating the concept of sample contrastive learning and bridging the gap between high-level and low-level tasks (Wu et al.
 2021; Chen et al.
@@
 2023).
-Image restoration tasks usually contain high-quality ground truth as learning target (positive sample), and more attention is paid on mining appropriate negative examples.
-For instance, Wu et al.
+Image restoration tasks usually contain high-quality ground truth as the learning target (positive sample), and more attention is paid to obtaining appropriate negative examples.
+For example, Wu et al.
 (Wu et al.
-2021) directly utilized the low-quality input as the negative sample and introduced the contrastive paradigm for the image dehazing task.
+2021) directly utilized low-quality input as the negative sample and introduced the contrastive paradigm for the image dehazing task.
 Wu et al.
@@
 Most recently, Zheng et al.
-proposed a progressively improved negative lower bound for image dehazing, employing multiple pre-trained models to provide consensual negatives (Zheng et al.
+utilized multiple pre-trained models to provide consensual negatives and proposed a progressively improved negative lower bound for image dehazing (Zheng et al.
 2023).
@@
 This provides a compact optimization space adaptively (pushing target model $f_{\theta}$ closer to assumed optimal $f_{\theta^{*}}$).
-While existing contrastive learning methods for low-level tasks, leveraging potent approaches like hard negative mining (Wu, Jiang, and Liu 2023) and curriculum learning strategies (Zheng et al.
+While existing contrastive learning methods for low-level tasks, leveraging potent approaches such as hard negative mining (Wu, Jiang, and Liu 2023) and curriculum learning strategies (Zheng et al.
 2023), have shown impressive performance, certain intrinsic limitations persist.
 Chiefly, there exists an over-reliance on task-oriented prior, leading to limited generalization capability across multiple image restoration tasks.
-Many existing methods exhibit a pronounced task-oriented bias, where negative sample generation is often influenced by prior knowledge and empirical evaluation centered around a singular target task (Wu, Jiang, and Liu 2023; Zheng et al.
+Many existing methods exhibit a pronounced task-oriented bias, where negative sample generation is often influenced by prior knowledge and empirical evaluation centered on a singular target task (Wu et al.
+2021; Wu, Jiang, and Liu 2023; Zheng et al.
 2023).
@@
 Taking this into consideration, we turn our focus to the target model itself, rather than sample selection.
-In recent literature, much of the spotlight has been directed towards negative sample collection, often overlooking a latent gem—the latency model—during the learning process.
+In recent literature, much of the spotlight has been directed towards negative sample collection, often overlooking a latent gem, the latency model, during the learning process.
 Specifically, the latency model, when operating within a close optimization step, shares strikingly similar parameters with the current model.
@@
 1.
-It becomes evident that, throughout the learning journey, the output of the latency model exhibits a sub-optimal but congruent distribution relative to the current anchor sample.
-This alignment offers the potential to derive ’hard’ negatives that are well-suited to the task at hand.
+It becomes evident that, throughout the learning journey, the output of the latency model exhibits a suboptimal but congruent distribution relative to the current anchor sample.
+This alignment offers the potential to derive ’hard’ negatives that are well suited to the task at hand.
 Furthermore, as the entire learning process is incrementally refined, the negatives in our model contrastive paradigm adopt a curriculum way naturally.
 Motivated by these insights, we put forth an innovative model contrastive paradigm for image restoration tasks.
-In contrast to the previous approaches with complex negative mining strategies, the core of our proposed model contrastive paradigm lies in the self-promoted negatives between latency and current models.
-Notably, it is task-agonist and general to diverse image restoration tasks.
-In detail, we proposed a model contrastive paradigm for various image restoration tasks (MCIR) with a Self-Prior guided Negative loss (SPN).
+In contrast to previous approaches with complex negative mining strategies, the core of our proposed model contrastive paradigm lies in the self-promoted negatives between latency and current models.
+Notably, it is a task-agonist and general to diverse image restoration tasks.
+In detail, we propose a model contrastive paradigm for various image restoration tasks (MCIR) with a Self-Prior guided Negative loss (SPN).
 Importantly, the proposed MCIR is compatible with existing approaches across different architectures and tasks.
@@
 2.
-One can find that retrained models achieve promising improvement across multiple tasks and architectures.
-More specifically, based on the lightweight EDSR (Lim et al.
+Retrained models can be found to achieve promising improvement across multiple tasks and architectures.
+More specifically, based on lightweight EDSR (Lim et al.
 2017) and SwinIR (Liang et al.
-2021), our retrained models display superior performance on Manga109 test dataset for ×$\times 4$ scale image super resolution, boasting improvements of 0.16 dB and 0.12 dB in term of PSNR.
+2021), our retrained models show superior performance on the Manga109 test dataset for ×$\times 4$ scale image super resolution, boasting improvements of 0.16 dB and 0.12 dB in terms of PSNR.
 Retrained IDT (Xiao et al.
 2022) achieves 0.38 dB and 0.7 dB gains on Rain200L and SPA datasets for image deraining.
-Moreover, for image dehazing, we gain a notable improvement of 3.41 dB and 0.57 dB compared to the original FFANet (Qin et al.
+Furthermore, for image dehazing, we gain a notable improvement of 3.41 dB and 0.57 dB compared to original FFANet (Qin et al.
 2020a) and DehazeFormer (Song et al.
@@
   • The paper proposes a novel approach MCIR through a task-agnostic model contrastive paradigm, which provides the adaptive generation of negative samples directly from the target model itself.
-Unlike conventional methods that manually apply negative samples to specific target task, the proposed model contrastive paradigm exhibits versatility across multiple tasks and models.
-  • The paper introduces the Self-Prior guided Negative loss (SPN) for image restoration, seamlessly compatible with existing methods.
+Unlike conventional methods that manually apply negative samples to a specific target task, the proposed model contrastive paradigm exhibits versatility across multiple tasks and models.
+  • The paper introduces the self-prior guided negative loss (SPN) for image restoration, which is seamlessly compatible with existing methods.
 SPN provides a simple to enhance existing image restoration models by integrating self-supervision principles within our model contrastive paradigm.
   • The paper demonstrates the effectiveness of the proposed approach by retraining existing models with the MCIR, which significantly improves image restoration across various tasks and architectures.
-Figure 2: Comparisons between models retrained by our proposed model contrastive paradigm and the originals.
-Retrained models can achieve remarkable improvements on various image restoration tasks.
-Details are provided in Section Experiment
+Figure 2: Comparisons between models retrained by our proposed model contrastive paradigm and the originals. Retrained models can achieve remarkable improvements on various image restoration tasks.
 Related Work
 Image Restoration
-Image restoration, aiming at recovering a high-quality image from the degraded one, is a fundamental problem in the fields of image processing and computer vision.
-It is indispensable in several domains, including photography, surveillance, medical imaging, and remote sensing (Su, Xu, and Yin 2022).
-In recent years, approaches based on deep learning have emerged as potent tools, providing significant advances for image restoration (Wang, Chen, and Hoi 2021a; Jiang et al.
+Image restoration, crucial in image processing and computer vision, aims to restore high-quality images from degraded versions (Su, Xu, and Yin 2022).
+Deep learning, particularly CNNs, has revolutionized this domain (Wang, Chen, and Hoi 2021a; Jiang et al.
 2023; Wang et al.
 2021a; Gui et al.
-2023b).
-Among these, Convolutional Neural Networks (CNN) have been utilized, resulting in a plethora of methods designed to address various image restoration tasks (Dong et al.
+2023b), with advancements in various CNN-based methods (Dong et al.
 2016; Zhang et al.
 2017; Zhang, Zuo, and Zhang 2018; Yang et al.
-2017a).
-Furthermore, remarkable strides have been made by the development of a spectrum of deep models, including proposals for residual or dense architectures (Lim et al.
-2017), the integration of attention mechanisms (Zhang et al.
+2017a), and the incorporation of novel architectures such as residual networks (Lim et al.
+2017), attention mechanisms (Zhang et al.
 2018; Liu et al.
-2019), and the implementation of UNet-based architectures (Ronneberger, Fischer, and Brox 2015; Jiang et al.
+2019), and UNet structures (Ronneberger, Fischer, and Brox 2015; Jiang et al.
 2020).
-In addition, some work explored the attention mechanism to boost the performance for image super-resolution (Zhang et al.
-2018; Liu et al.
-2018; Mei, Fan, and Zhou 2021), image deraining (Li et al.
-2018a; Wang et al.
-2019a), image deblurring (Suin, Purohit, and Rajagopalan 2020), and so on.
-On the frontier of recent development, Transformer architectures achieves significant breakthroughs in computer vision (Liu et al.
-2021; Dosovitskiy et al.
-2021).
-This trend has spurred the development of Transformer-based interventions for image restoration tasks such as image super-resolution (Liang et al.
-2021), image dehazing (Song et al.
-2023; Zamir et al.
-2022), image deraining (Xiao et al.
-2022), image deblurring (Wang et al.
-2022), and low-light image enhancement (Wang et al.
-2023).
-We are witnessing a surge in the performance of increasingly complex architectures and there is a growing demand for applications in real-world systems.
-Many researcher have turned attention to lightweight methods (Jiang et al.
+As well as some lightweight architectures were proposed for practical application (Jiang et al.
 2021; Lamba and Mitra 2021; Hui et al.
@@
 2023a).
+Recent trends include Transformer-based architectures (Liu et al.
+2021; Dosovitskiy et al.
+2021), delivering breakthroughs in many image restoration tasks (Liang et al.
+2021; Song et al.
+2023; Zamir et al.
+2022; Xiao et al.
+2022; Wang et al.
+2022, 2023).
+In this work, we turn our attention to the optimizing strategy and exploit an effective model contrastive learning method to refresh the performance existing models.
 Contrastive Learning
-Self-supervised learning has achieved great success in high-level vision tasks, especially those contrastive learning methods (Gui et al.
+Self-supervised learning, particularly contrastive methods, has seen remarkable success in high-level vision tasks (Gui et al.
 2023a).
-For low-level image restoration tasks, many researchers attempted to bridge the gap between low-level and high-level tasks and exploit more self-supervised information to promote image restoration tasks (Wang et al.
+In the field of low-level image restoration, researchers are exploring the integration of self-supervised regularization to bridge the gap between low- and high-level tasks (Wang et al.
 2021b; Wu et al.
@@
 2023; Wu, Jiang, and Liu 2023).
-In contrast to unsupervised learning for high-level vision tasks, low-level tasks usually have high-quality ground-truth images.
-Discovering informative negative samples is crucial for low-level tasks.
+Typically, low-level tasks often have high-quality ground truth, making the identification of informative negative samples vital.
 Wu et al.
-took the low-quality input as the negative sample and proposed a perceptual-based contrastive loss to achieve a more compact convergence space (Wu et al.
+employed low-quality inputs as negative samples and introduced a perceptual-based contrastive loss for more effective convergence (Wu et al.
 2021).
 Wu et al.
-proposed a practical contrastive learning framework for super-resolution single image and proposed the hard negative construction (Wu, Jiang, and Liu 2023).
-By interpolating the negative information, they further improve the performance for target image restoration task.
-However, adaptation and generalization of these methods to various image restoration models and tasks are limited.
-Compared to existing methods (Wu et al.
+proposed a practical contrastive learning framework for single image super-resolution tasks, enhancing performance through hard negative construction and negative information interpolation (Wu, Jiang, and Liu 2023).
+Nonetheless, the adaptability and generalization of these methods across diverse image restoration models and tasks remain a challenge.
+Our work distinguishes itself from existing methods (Wu et al.
 2021; Wu, Jiang, and Liu 2023; Zheng et al.
-2023), our work presents a task-agnostic model contrastive paradigm for image restoration, utilizing a self-prior guided negative loss.
-This approach simplifies negative sample generation and offers a versatile framework for various image restoration tasks, setting it apart from existing methods.
+2023) by introducing a task-agnostic model contrastive paradigm for image restoration.
+This novel approach shifts the focus from negative samples to negative models, offering a straightforward yet effective framework for a variety of image restoration tasks.
 Method
-Figure 3: Comparisons of different learning paradigms for image restoration tasks.
-(a) Depicts the vanilla end-to-end learning paradigm, where target image restoration models are trained with supervised reconstruction loss; (b) Illustrates the sample contrastive paradigm adopted in existing methods, leveraging additional information from manually defined negative samples; (c) Showcases our model contrastive paradigm, which generates negative samples from the latency target model, providing a task-agnostic and adaptive approach to various image restoration tasks.
 Overall Framework
 Deep learning-based image restoration methodologies have recently made considerable breakthroughs across multiple restoration tasks, including image super-resolution, image dehazing, image deblurring, image deraining, and so on.
-In this paper, we focus our attention on the vanilla end-to-end learning process in deep learning methods.
-To facilitate clarity, we provide a succinct illustration of the end-to-end image restoration framework, visualized in Fig.
-3 (a).
-Given a low-quality input image denoted as $I^{LQ}$, the target model $f_{\theta}$, where $\theta$ represents model parameters, processes low-quality input to produce a reconstructed image, $I^{Rec}$.
-Concurrently, its high-quality counterpart is denoted by $I^{HQ}$ and the optimizing of the target model is steered by the minimization of the reconstruction loss $\mathcal{L}_{rec}$.
-Delving further, Fig.
-3 (b) showcases the contrastive learning strategies tailored for image restoration tasks.
-Typically, researchers harness the power of negative samples, exploiting informative task-prior specific negatives and extend the vanilla end-to-end learning paradigm by the negative constraint $\mathcal{L}_{neg}$ (Wu et al.
+In this paper, we focus our attention on the vanilla end-to-end framework with deep learning methods.
+Given a low-quality input image $I^{LQ}$, the target model $f_{\theta}$ (where $\theta$ denotes the model parameters) processes this input to produce a reconstructed image $I^{Rec}$.
+The high-quality counterpart of the image is represented as $I^{HQ}$.
+The optimization of $f_{\theta}$ is guided by minimizing the reconstruction loss $\mathcal{L}_{rec}$, defined as:
+$\min_{\theta}\mathcal{L}_{rec}(f_{\theta}(I^{LQ}),I^{HQ}),$ (1)
+where $\mathcal{L}_{rec}$ typically depends on metrics such as Mean Absolute Error (MAE) and Mean Squared Error (MSE).
+Delving further, existing contrastive learning methods for image restoration (Wu et al.
 2021; Zheng et al.
-2023; Wu, Jiang, and Liu 2023).
-In this paper, we propose an innovative model contrastive paradigm for image restoration tasks, shown in Fig.
-3 (c).
-Instead of utilizing task-specific and predefined negative samples, we craft our negatives directly from the target model itself.
-At this point, we denote the target model $f_{\theta}$, indicative of the optimized one at the current training iteration $t$.
-Subsequently, we employ a distinct step, $s$, to establish a latency model $f_{\theta^{{}^{\prime}}}$, also referred to as the negative model, to generate adaptive negative samples for the training of the target model.
-The overall pipeline is straightforward, and the detailed implementation of negative models and associated loss functions will be discussed in the following sections.
+2023; Wu, Jiang, and Liu 2023) utilize negative samples $I^{Neg}$, extending the standard end-to-end learning with a negative loss $\mathcal{L}_{neg}$:
+$\min_{\theta}\mathcal{L}_{rec}(f_{\theta}(I^{LQ}),I^{HQ})+\mathcal{L}_{neg}(f_{\theta}(I^{LQ}),I^{Neg}),$ (2)
+where the negative samples are predefined and generated based on task-specific priors, which limit their generalization capability across diverse image restoration tasks.
+In this work, we introduce a novel model contrastive paradigm for image restoration, shifting from task-specific negative samples to negatives derived from the target model itself.
+We designate the target model as $f_{\theta}$, representing its state in a given training iteration $t$.
+A key innovation is the development of a latency model, $f_{\theta^{{}^{\prime}}}$ (the negative model), which is utilized to produce adaptive negative samples for training $f_{\theta}$.
+This approach simplifies the overall process in existing methods and provides a general framework for image restoration.
+The specifics of implementing negative models and the corresponding loss functions will be elaborated in subsequent sections.
 Learning from History
 We propose a practical implementation for leveraging latency models to generate informative negative samples during training, as introduced earlier.
-This method circumvents the challenges of significant gaps between model checkpoints and the impracticality of frequent processing, especially with larger models.
+This method avoids the challenges of significant gaps between model checkpoints and the impracticality of frequent processing, especially with larger models.
 In detail, we introduce the efficient strategy of exponential moving averages (EMA) to achieve a smooth negative model.
-The update equation for the negative model is:
-$\begin{split}&\theta^{{}^{\prime}}=w\theta^{{}^{\prime}}+(1-w)\theta,\\ &\mathrm{s.t.}~{}t\%s=0.\end{split}$ (1)
+The update equation for the negative model is as follows:
+$\theta^{{}^{\prime}}=w\theta^{{}^{\prime}}+(1-w)\theta,~{}\mathrm{s.t.}~{}t\%s=0.$ (3)
 In this schema, $w$ denotes the update weight, $s$ is the update step, and $t$ captures the current iteration.
-To ensure the preservation of latency parameters, we adopt a long step $s$ and selectively update the negative model $f_{\theta}$ at intervals of every $s$ iterations.
+To ensure the preservation of latency parameters, we adopt a long step $s$ and selectively update the negative model $f_{\theta}$ at intervals of every $s$ iteration.
 Self-Prior Guided Negative Loss
 At the heart of our model contrastive paradigm, a loss function mediates between the target reconstruction $I^{Rec}$ and its negative counterpart, $I^{Neg}$.
-Herein, we take the pre-trained VGG (Simonyan and Zisserman 2015) network as the embedding network to map the samples into a latent feature space, where $f^{Rec}=\text{VGG}(I^{Rec})$ and $f^{Neg}=\text{VGG}(I^{Neg})$.
+Here, we take the pre-trained VGG (Simonyan and Zisserman 2015) network as the embedding network to map the samples into a latent feature space, where $f^{Rec}=\text{VGG}(I^{Rec})$ and $f^{Neg}=\text{VGG}(I^{Neg})$.
 Then the proposed negative loss $\mathcal{L}_{neg}$ is formulated as
@@
 $\mathcal{L}^{N}_{neg}=\frac{1}{N}\sum_{i=1}^{N}{\|f^{Rec}-f^{Neg}_{i}\|}_{1},$ (5)
-where $N$ denotes the total number of negative models and $f_{i}^{Neg}$ corresponds to the latent feature of the $i$th negative sample.
+where $N$ denotes the total number of negative models and $f_{i}^{Neg}$ corresponds to the latent feature of the $i$-th negative sample.
 Compared to existing methods, our negative regularization is a self-prior guided loss function, where the negative samples stem from the target model itself.
@@
 To validate the prowess of our model-based contrastive paradigm by $\mathcal{L}_{neg}$, we have retrained numerous exiting methods with it, testing across different image restoration tasks and architectures.
-The precise formulation of the reconstruction loss $\mathcal{L}_{rec}$ is contingent on the method under retraining.
+The precise formulation of the reconstruction loss $\mathcal{L}_{rec}$ is dependent on the retraining method.
 Generally, the total loss function within our model-based contrastive paradigm is defined as:
@@
 This simple formulation allows us to incorporate the proposed SPN with existing image restoration methods, enhancing their flexibility and adaptability to various tasks.
-Method Architecture Scale
+Methods Architecture Scale
 Avg.
@@
 PSNR/SSIM
-EDSR-light CNN ×$\times 2$ 32.06/0.9303 33.57/0.9175 32.16/0.8994 31.98/0.9272 30.54/0.9769
-+SPN (Ours) 32.19/0.9313 33.67/0.9182 32.21/0.9001 32.23/0.9297 30.64/0.9772
+EDSR-light CNN ×$\times 2$ 34.06/0.9303 33.57/0.9175 32.16/0.8994 31.98/0.9272 38.54/0.9769
++SPN (Ours) 34.19/0.9313 33.67/0.9182 32.21/0.9001 32.23/0.9297 38.64/0.9772
 EDSR-light ×$\times 4$ 28.14/0.8021 28.58/0.7813 27.57/0.7357 26.04/0.7849 30.35/0.9067
@@
 Experimental Settings
+In this section, we apply the following four image restoration tasks to test and validate the effectiveness of the proposed method.
 Image Super-Resolution
@@
 Image Deblurring
-For image deblurring, we retrain and evaluate NAFNet (Chen et al. 2022a) on GoPro dataset.
+For image deblurring, we retrain and evaluate NAFNet (Chen et al. 2022a) on GoPro dataset (Nah, Hyun Kim, and Mu Lee 2017).
 Image Deraining
@@
 2017b), DID-Data (Li et al.
-2018b), DDN-Data (Fu et al.
+2018), DDN-Data (Fu et al.
 2017), and SPA-Data (Wang et al.
-2019b).
+2019).
 We adopt the Transformer-based IDT model (Xiao et al.
@@
 Comparison Results
-Method
+Methods
 Avg.
 PSNR/SSIM
-Rain100L
+Rain200L
 PSNR/SSIM
-Rain100H
+Rain200H
 PSNR/SSIM
@@
 Image Super-Resolution
-We first investigate the image super-resolution task, utilizing both CNN-based EDSR and Transformer-based SwinIR as baselines to evaluate our proposed self-pprior guided negative regularization.
+We first investigate the image super-resolution task, utilizing both CNN-based EDSR and Transformer-based SwinIR as baselines to evaluate our proposed self-prior guided negative regularization.
 Results presented in Tab.
@@
 Our proposed model contrastive paradigm can achieve considerable improvements when retraining lightweight models such as EDSR and SwinIR, achieving average gains of 0.07 dB and 0.09 dB on scale $\times$4, respectively.
-Moreover, the retrained EDSR model demonstrats a noteworthy average improvement of 1.3 dB for the scale $\times$2 task.
-Notably, there are substantial improvements by our retrained on the Urban100 and Manga109 datasets.
+Moreover, the retrained EDSR model demonstrates a noteworthy average improvement of 1.3 dB for the scale $\times$2 task.
+In particular, there are substantial improvements from our re-training on the Urban100 and Manga109 datasets.
 This implies that our model contrastive paradigm facilitates a more accurate and compact convergence in the restoration process.
@@
 Our proposed approach exhibits its general applicability to larger models, providing an average improvement of 0.05 dB.
-It is worth noting that the improvements seen in larger models were relatively smaller when compared to the lightweight models.
+It is worth noting that the improvements seen in the larger models were relatively smaller compared to the lightweight models.
 This observation is consistent with the understanding that lightweight models, owing to their smaller capacity, may struggle to find an optimal solution.
@@
 Visual results are showcased in Fig.
-4.
+3.
 It is also evident that our model contrastive paradigm can improve existing methods, particularly in bringing more precise textures.
 Image Deraining
-We take the state-of-the-art IDT (Xiao et al.
-2022) as the baseline and retrain it by the proposed model contrastive paradigm on five datasets ,and Tab.
+We take the advanced IDT (Xiao et al.
+2022) as a baseline and retrain it by the proposed model contrastive paradigm on five datasets, and Tab.
 2 provides a comprehensive comparison.
-The retrained IDT, achieves an average PSNR improvement of 0.26 dB across all test datasets.
-For instance, it showcase a gain of 0.7 dB in comparison to its original on SPA dataset.
+The retrained IDT achieves an average PSNR improvement of 0.26 dB across all test datasets.
+For instance, it showcases a gain of 0.7 dB in comparison to its original in the SPA dataset.
 Furthermore, visual results are illustrated in Figure 4, accompanied by a divergence map.
-It intuitively highlights the enhancements in the degraded regions.
-These insights underscore the capability of our model contrastive paradigm to refine the target degradation.
+It intuitively highlights enhancements in the degraded regions.
+These insights underscore the ability of our model contrastive paradigm to refine target degradation.
+Table 3: Comparison results of image dehazing. The results of our retrained models are in bold
 Methods ITS RESIDE-6K
@@
 (Ours) DehazeFormer-B+SPN 38.41 0.994 31.57 0.981
-Table 3: Comparison results of image dehazing. Reults of our retrained models are in bold.
-Figure 6: Visual results of FFANet and our retrained one for image dehazing.
-Method MIMO-UNet HINet MAXIM Restormer UFormer NAFNet NAFNet+SPN (Ours)
-PSNR 32.68 32.71 32.86 32.92 32.97 32.87 32.93 (+0.06)
-SSIM 0.959 0.959 0.961 0.961 0.967 0.9606 0.9619 (+0.0013)
-Table 4: Comparison results of image deblurring. We take NAFNet (Chen et al. 2022a) as the benchmark and retrain it with the proposed model contrastive paradigm on GoPro dataset.
+Table 3: Comparison results of image dehazing. The results of our retrained models are in bold
 Image Dehazing
@@
 2020b) and DehazeFormer (Song et al.
-2023), with the outcomes detailed in Fig.
+2023), with the results detailed in Fig.
 3.
 From the table, a remarkable enhancement is achieved over FFANet, exhibiting gains of 3.41 dB and 0.69 dB on the indoor and mixed test datasets, respectively.
-Concurrently, the Transformer-based model IDT also record advancements, evident across various model scales.
+Concurrently, the Transformer-based model DehazeFormer also records advancements, evident across various model scales.
 Specifically, the retrained DehazeFormer-B achieves the most pronounced improvement of 0.57 dB on the indoor test dataset.
-Intriguingly, the retrained FFANet even outperformes the DehazeFormer.
+Intriguingly, the retrained FFANet even outperforms the DehazeFormer.
 It is reasonable that FFANet is a CNN-based model, which has a large model capacity with heavy parameter counts, and our model contrastive paradigm appears to steer FFANet towards more optimized results.
 Some visual results are presented in Fig.
-6.
-One can find that retrained FFANet is clearer with less artifacts.
+5.
+One can find that the retrained FFANet is clearer with fewer artifacts.
 Image Debluring
@@
 2022).
-Methods Task & Dataset PSNR SSIM
-FFANet (Qin et al. 2020b) 36.39 0.9886
-+CR (Wu et al. 2021) Image Dehazing 36.74 0.9906
-+CCR (Zheng et al. 2023) (SOTS-indoor ) 39.24 0.9937
-+SPN (Ours) 39.80 0.9947
-EDSR (Lim et al. 2017) SISR 26.04 0.7849
-+PCL (Wu, Jiang, and Liu 2023) (Urban100) 26.07 0.7863
-+SPN (Ours) 26.12 0.7878
-Table 5: A comparative analysis of the existing contrastive paradigms versus our proposed model contrastive approach.
-Existing methods are task-oriented and proposed for image dehazing and image super-resolution separately.
-Our model contrastive paradigm is task-agonist and general to improve diverse image restoration tasks.
+Figure 5: Visual results of FFANet and our retrained one for image dehazing.
 Comparison to Existing Sample Contrastive Paradigm
@@
 We furnish a comprehensive comparison between exiting contrastive approaches and our innovative model contrastive paradigm.
-For image dehazing task, we examine previous methods based on FFANet, including contrastive regularization (CR) (Wu et al.
+For the image dehazing task, we examine previous methods based on FFANet, including contrastive regularization (CR) (Wu et al.
 2021), curricular contrastive regularization (CCR) (Zheng et al.
@@
 Instead of merely employing low-quality input images as negative samples, CCR achieves significant advancements by utilizing consensus negatives from pre-trained models and the curriculum learning approach.
-Intriguingly, our SPN outperforms other methods, achieving the best improvement without any additional prior knowledge.
+Intriguingly, our SPN outperforms other methods, achieving the best improvement without additional prior knowledge.
 In addition, for image super-resolution (Wu, Jiang, and Liu 2023), our model contrastive paradigm again registers the highest gains, as presented in Tab.
@@
 This consistent performance underscores the fact that our proposed model contrastive paradigm is task-agnostic and highly effective for diverse image restoration tasks.
+Methods MIMO-UNet HINet MAXIM Restormer UFormer NAFNet NAFNet+SPN (Ours)
+PSNR 32.68 32.71 32.86 32.92 32.97 32.87 32.93
+SSIM 0.959 0.959 0.961 0.961 0.967 0.9606 0.9619
+Table 4: Comparison results of image deblurring. We take NAFNet (Chen et al. 2022a) as the benchmark and retrain it with the proposed model contrastive paradigm on GoPro dataset.
+Methods Task & Dataset PSNR SSIM
+FFANet 36.39 0.9886
++CR Image Dehazing 36.74 0.9906
++CCR (SOTS-indoor ) 39.24 0.9937
++SPN (Ours) 39.80 0.9947
+EDSR SISR 26.04 0.7849
++PCL (Urban100) 26.07 0.7863
++SPN (Ours) 26.12 0.7878
+Table 5: A comparative analysis of the existing contrastive paradigms versus our proposed model contrastive approach.
+Typically, existing methods are task-oriented and are proposed for image dehazing (CR (Wu et al.
+2021) and CCR (Zheng et al.
+2023)) and image super-resolution (PCL (Wu, Jiang, and Liu 2023)), separately.
+Our model contrastive paradigm is task-agonist and outperform existing methods.
 Negative Model - Random Pre-trained Default
@@
 Impact of Negative Model
-To provide a deeper understanding for the role of our negative model, our ablation study involves the exploration of different configurations of the negative model, including an randomly initialized and fixed SwinIR model, a pre-trained and fixed SwinIR model, and our default one.
-The results of these different configurations, which shed light on the efficacy of our proposed approach, are presented in Tab.
-6.
-’-’ presents the baseline model without contrastive negative model.
-Firstly, negative model with random and fixed parameters can not afford effective negatives and impedes the target SR model.
-When we set the negative model as the pre-trained and fixed one, it enhances the target model but achieves worse performance than ours.
-We think the adaptive and curricular negatives generated by model contrastive paradigm is essential to provide a smooth negative regularization, which is consistent with some existing methods.
+This ablation study aims to elucidate the role of our negative model by investigating various configurations, including a randomly initialized and fixed SwinIR model, a pre-trained and fixed SwinIR model, and our standard model.
+The results, detailed in Tab.
+6, highlight the effectiveness of our approach.
+The baseline model is denoted by ’-,’ indicating the origin SwinIR.
+We found that a negative model with random and fixed parameters hinders the target SR model, failing to provide effective negatives.
+Conversely, using a pre-trained and fixed negative model shows improvement, but not as much as our model contrastive.
+This suggests the importance of adaptive and curriculum-based negative generation in our model contrastive paradigm, aligned with the findings of existing methods (Zheng et al.
+2023).
 Step $s$ 100 500 1000 2000 All
@@
 Impact of Negative Step
-We conducted an ablation study, where FFANet was retrained with various negative models, each differentiated by its updating iteration, $s$.
+We conducted an ablation study, in which FFANet was retrained with various negative models, each differentiated by its update iteration, $s$.
 The results are tabulated in Tab.
@@
 (3), and results are presented in Tab.
-8, where 0 value means the original $L_{1}$ loss without our negative regularization.
-When $\lambda$ set as 1e-2 or 1e-3, there is collapse during training, especially in the early stage.
-This is reasonable that large negative loss can influence the model updating towards to the optimal direction.
+8, where the value 0 means the original loss $L_{1}$ without our negative regularization.
+When $\lambda$ is set to 1e-2 or 1e-3, there is collapse during training, especially in the early stage.
+This is reasonable because large negative loss can influence the model updating towards the optimal direction.
 When $\lambda$ is 1e-4, our model contrastive paradigm achieves the best performance.
-Considering that the proposed model contrastive paradigm is task-agonist and adaptive to various image restoration tasks, in this paper, we take $\lambda=1e-4$ as the default value across different tasks.
-$\lambda$ Value 0 1e-2 1e-3 5e-4 1e-4 1e-5
+Considering that the proposed model contrastive paradigm is general to various image restoration tasks, in this paper we take $\lambda=1e-4$ as the default value across different tasks.
+$\lambda$ 0 1e-2 1e-3 5e-4 1e-4 1e-5
 EDSR-light 28.58 - - 28.56 28.63 28.60
@@
 As $w$ increases, it becomes evident that the optimal performance is attained at $w=0.1$.
-Based on our empirical observations, we consistently utilize $w=0.1$ as the default setting in all subsequent experiments.
+On the basis of our empirical observations, we consistently use $w=0.1$ as the default setting in all subsequent experiments.
 Discussion and Limitation
-In this paper, we proposed a task-agnostic model contrastive paradigm for low-level tasks, improving existing models.
-While our experiments cover a range of tasks, areas like JPEG artifact removal, image denoising, and certain real-world scenarios remain unexplored.
-Our ablation studies focused on image super-resolution, as shown in Tables 6 and 7, but we did not extensively evaluate our hyper-parameters across all tasks.
+In this paper, we proposed a novel model contrastive paradigm for low-level image restoration tasks, improving existing models.
+Although our experiments cover a range of tasks, areas such as JPEG artifact removal, image denoising, and certain real-world scenarios remain unexplored.
+Our ablation studies focused on image super-resolution, as shown in Tab.
+6 and  7 but we did not extensively evaluate our hyperparameters across all tasks.
 These unexamined aspects offer valuable opportunities for future research.
 Conclusion
-In this study, we propose innovative model contrastive paradigm for various low-level tasks.
-Compared to the task-oriented negatives in existing methods, the proposed model contrastive paradigm, constructing negatives from the target model itself, a is task-agonist, general to various image restoration tasks by a Self-Prior guided Negative loss (SPN).
-Our proposed SPN is straightforward to implement and does not rely on any task or data prior.
+In this study, we propose an innovative model contrastive paradigm for various low-level tasks.
+Compared to the task-oriented negatives in existing methods, the proposed model contrastive paradigm, constructing negatives from the target model itself, is task-agonist and general to various image restoration tasks by a self-prior guided negative loss (SPN).
+Our proposed SPN is straightforward to implement.
 We have retrained several image restoration models, and they achieve significant improvements across various tasks and architectures.
 In the future, we believe it would be meaningful to evaluate our proposed paradigm in more dense prediction tasks, potentially offering fresh insights and advances for the community.
-Acknowledgments The research was supported by the National Natural Science Foundation of China (U23B2009, 92270116).
+Acknowledgments
+The research was supported by the National Natural Science Foundation of China (U23B2009, 92270116).